# Breaking the Frame: Visual Place Recognition by Overlap Prediction

### v2

@@ 4 Experimental Results
Our experiments include estimating two-view epipolar geometry and localizing query images from retrieved images. The VOP model is trained on 153 scenes from the MegaDepth dataset [28], which includes Internet-sourced images and depths reconstructed via Structure-from-Motion (SfM) and multi-view stereo techniques. Beyond the MegaDepth testing scenes, we evaluate the generalization ability of the proposed VOP on 11 scenes from PhotoTourism [49], 13 scenes from ETH3D [45], and the InLoc dataset [51]. We benchmark the proposed method against several competitors, including the DINOv2, CNN-based retrieval methods like NetVLAD [2] and CosPlace [6], as well as the state-of-the-art (SOTA) retrieval method AnyLoc [25] and DINOv2-SALAD [24] utilizing the same DINOv2 backbone as VOP. Besides those global-only methods, we compare VOP with the SOTA reranking-based methods: the ViT-based $R^{2}$ormer [62] and the CNN-based Patch-NetVLAD [20]. Also, we finetune Cosplace* on the same dataset our method was trained on to demonstrate that the improved performance is not due to the different training data. In addition, we test reranking methods [20] [62] on the shortlists created by DINOv2 [CLS] token, similarly as done for VOP, marked with $\dagger$ in the tables.
@@
Training Details. The models underwent training on the 153 MegaDepth scenes, employing contrastive loss to differentiate negative and positive patch pairs. Each scene contributes 150 positive and 150 negative image pairs to the training set, standardized to an image resolution of $224\times 224$ pixels. We use the DINOv2 pre-trained model with a patch size of $14$ to extract $1024$-dimensional descriptors and reduce them to 256 dimensions by a fully-connected layer, i.e., the first layer or our training encoder head. Negative images, characterized by zero overlaps (e.g., sourced from distinct scenes), contrast with positive samples randomly selected from the pairs with 10% to 70% overlapping patches. Notably, the contrastive loss framework is predicated on patch-level negative/positive distinctions rather than entire images. Note that cosine similarity is used in the contrastive loss. Following LightGlue [29], patch-level labels are generated from depth information and visual features, utilizing the GlueFactory library [36].
The training was conducted over 30 epochs with a batch size of 64, employing the validation losses as the primary criterion for best checkpoint selection and utilizing a learning rate of $41e-$ 4. The voting scheme and the radius for neighbor searching were optimized on the validation set by AUC scores, which included a single scene for this purpose. During inference, scenes “0015” (comprising approximately 0.4K images) and “0022” (encompassing approximately 0.5K images) from MegaDepth were designated for testing, enabling a focused evaluation of the model performance on these specific subsets.
@@
Evaluation Metrics. In the evaluation, we assess the relative poses between the query and the top-$k$ retrieved images using robust estimation techniques. This involves deploying SuperPoint [13] and the state-of-the-art LightGlue [29, 44] for feature detection and matching, followed by OpenCV-RANSAC [17] running for 10K iterations to compute the relative poses. The performance is measured against ground truth (GT) poses by calculating the Area Under the Recall curve (AUC) thresholded at $10^{\circ}$ [59, 8], along with the median pose error. The RANSAC threshold is optimized from a set of pixel values $\{0.5,1,2,3\}$ on the training set. Also, we report the average number of inliers across the retrieved image pairs found by RANSAC. Note that the number of inliers is not an indicator of the accuracy of the estimated pose, as discussed and verified in MQNet [4] and FSNet [5]. In scenarios where $k>1$, for each query, we follow the practical approach and select only the image pair that has the highest number of inliers. This comprehensive evaluation allows us to thoroughly assess the performance not just in retrieving relevant images but also in accurately estimating their relative poses.
@@
Additionally, we visualize some retrieval examples with their baselines and accuracy in Fig. 8 to show the aim of designing retrieval methods tailored for geometric challenges, e.g., pose estimation.
4.1 Pose Estimation on MegaDepth
Table 1 presents the results on the MegaDepth, highlighting the performance of the proposed VOP. The proposed method secures comparable scores for the AUC@10 criterion and exhibits the lowest median pose errors among the SOTA baselines. The underlined results mark the optimal configuration for relative pose accuracy, evidencing the superior performance of the proposed VOP approach when utilizing the top-5 images. This configuration outperforms the global-only retrieval baselines (the first seven rows), reinforcing the significance of selecting correct metrics and model configurations tailored to the specific demands of real-world downstream applications. Our method outperforms the recent reranking approaches $R^{2}$ormer [62] and Patch-NetVLAD [20] on MegaDepth when using their pretrained models. When using the same shortlist we use, their performance (with $\dagger$) is marginally better than ours. The fine-tuned CosPlace (CosPlace*) achieves better accuracy than the pretrained model, especially achieving good results on top-10 pairs. Next, we will show the generalization and robustness of the compared methods on unseen data.
@@ 4.1 Pose Estimation on MegaDepth
VOP 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
Table 1: Relative pose estimation on MegaDepth using top-$k$ images retrieved by different methods, with the best results in bold at each $k$ and the overall best underlined. results are tested using reranking methods prefiltered by the [CLS] tokens from DINOv2.
@@ 4.2 Generalization Experiments
AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
NetVLAD [2] 89.5 90.3 89.9 0.74 0.74 0.79 365.8 147.3 440.1
CosPlace [6] 88.3 89.7 89.7 0.77 0.77 0.78 349.7 426.8 437.2
CosPlace* [6] 43.6 79.7 85.9 38.3 0.79 0.76 162.6 279.0 360.5
DINOv2 [35] 87.1 89.2 88.9 0.69 0.67 0.67 338.9 378.2 399.8
AnyLoc [25] 88.9 90.2 90.2 0.76 0.78 0.77 373.2 425.1 438.0
SALAD [24] 89.6 90.2 89.9 0.73 0.76 0.78 385.9 411.2 442.0
P-NetVLAD [20] 86.7 88.4 90.2 0.81 0.83 0.81 333.3 412.4 466.8
P-NetVLAD [20] 80.5 87.7 89.8 0.87 0.65 0.66 273.2 319.1 355.3
$R^{2}$ormer[62] 49.1 79.8 90.1 35.2 6.94 0.83 116.5 252.8 364.9
$R^{2}$ormer[62] 50.6 81.5 88.2 27.3 0.87 0.80 137.0 318.3 401.5
VOP 83.6 91.1 90.4 0.69 0.61 0.72 293.9 414.7 460.0
Table 2: Relative pose estimation on the ETH3D dataset [45] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. The proposed VOP is trained on the MegaDepth dataset. results are tested using reranking methods prefiltered by the [CLS] tokens.
@@
VOP 62.5 67.5 65.4 2.63 2.00 2.32 224.5 286.1 281.3
Table 3: Relative pose estimation on the PhotoTourism dataset [49] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. VOP is trained on the MegaDepth dataset. * shows fine-tuned Cosplace on Megadepth.
Methods
@@
Avg. Accuracy (%) $\uparrow$ 70.1 69.6 56.4 70.5 73.2 72.3 66.2 65.3 75.1
Avg. med. pose error ($\downarrow$ 12.1 12.0 12.0 11.9 12.0 12.2 11.9 11.8 11.5
Table 5: Average accuracy (in percentage) and median pose error (in degrees) on the retrieved top-5 images on the Megadepth [28], ETH3D [45], PhotoTourism [49], and InLoc [51] datasets. InLoc is excluded from the median error, as the official website does not report it. indicates prefiltering by the [CLS] tokens. * shows fine-tuned Cosplace on Megadepth. The best results are shown in bold.
InLoc. Additionally, we ran all methods on the InLoc dataset [51] (without fine-tuning) using hloc [44] for localization to show their generalization capabilities in the case of a large domain gap (in/outdoor). Following hloc [44], we retrieve top-40 database images for all the retrieval methods. For the reranking ones (including VOP), we rerank the top-100 shortlist created by [CLS] token. Table 4 shows the proposed VOP achieves the second highest accuracy, showcasing its generalization abilities. Note that we only show results for P-NetVLAD and $R^{2}$ormer with DINOv2 [CLS] token as their original version fails on this dataset.
In summary, we show the average accuracy and median pose errors averaged over all tested datasets in Table 5. Note that InLoc is excluded from the median pose error, as the official website does not report it. Overall, the proposed VOP leads to the highest accuracy and lowest pose errors. Fine-tuning Cosplace on Megadepth reduces its accuracy on average, demonstrating that our improved performance is not simply due to the different training data. Again, we do not report results for P-NetVLAD and $R^{2}$ormer without DINOv2 [CLS] token due to their failure on InLoc.
Figure 6: AUC@10 scores on the queries with top-5 images from database images lower than the limits along the horizontal axis.
@@ 4.3 Discussions
Storage & Query Speed. As we reduce the dimensionality of the DINOv2 features to 256, the embeddings of all patches of each image need a total of 512 kB, while, for AnyLoc, the storage per image is 384 kB. While we require slightly more storage than AnyLoc [25], the difference is small. Compared to storing the local features in the reranking-based methods, VOP costs less. In addition, we compare the time of querying an image from the database of different sizes using AnyLoc or VOP in seconds. Prefiltering top 20 images by DINOv2 [CLS] token and running VOP for reranking to get top-1 out of 500 images cost 0.009 seconds, while 0.003 for AnyLoc. Querying top-1 from 5K images (prefiltered to 100), our method costs 0.03 seconds, while AnyLoc runs in 0.02s. Note that VOP needs much more time for radius search without prefiltering. We recommend using VOP as a reranking method in retrieval.
Image Patching. We investigated the number of patches to be used to split the images. The experiments were conducted on the testing scenes of MegaDepth, with all the images resized to $224\times 224$. From the trained VOP model with a patch size of $14^{2}$, we can extract 256 patch descriptors. Then, average pooling is applied to aggregate the patch descriptors to different patch sizes, such as $28^{2}$, $56^{2}$, $112^{2}$, and $224^{2}$. For example, patch size = 224 will lead to a single patch of an image. The retrieval is done on the same prefiltered image list as discussed. As shown in Table 6, the aggregated patches e.g., patch size=$224^{2}$, perform worse than $14^{2}$ on MegaDepth pose estimation and could not generalize well on Inloc localization. This demonstrates that patch-level features can potentially improve estimated pose and other geometric problems.
MegaDepth - top5 DUC1 DUC2
$224^{2}$ 67.0 2.09 30.8 24.4
$112^{2}$ 67.0 2.09 38.4 38.9
$56^{2}$ 66.5 2.17 48.5 57.3
$28^{2}$ 65.9 2.29 59.1 72.5
$14^{2}$ 67.6 2.03 72.2 77.1
Table 6: Ablations on different patch sizes used in inference time. We show the AUC@10 scores and median pose errors of the top-5 retrieved images on MegaDepth [49], and the recall@, 0.5m on the localization data, Inloc [51] (DUC1, DUC2).
Qualitative Results. Most VPR methods prioritize retrieving similar images, typically resulting in short baselines that are not suitable for reconstruction. These goals conflict: the most similar images often produce short baselines, making pose estimation unstable. We aim to move beyond traditional similarity metrics and design retrieval methods tailored for geometric challenges, such as selecting images suitable for pose estimation. We visualize three query examples in Fig. 8 with their top-1 retrieved images using different methods. VOP results in low pose errors as we find images with reasonable baselines for stable pose estimation.
5 Ablations
We investigate the influence of image patches, different patch sizes, the threshold used in radius search and the influence of the data augmentation on test accuracy and generalization capability.
Radius Search Threshold. One of the parameters of the proposed VOP is the radius search threshold when selecting similar patches in the embedding space. To understand how sensitive VOP is to the setting of this threshold, we show tuning results on the validation set of the MegaDepth dataset. We show the results with different thresholds (horizontal axis) on the validation set in Fig. 7. While the accuracy varies with the threshold, setting it on the validation set and using this fixed threshold on all tested large-scale datasets leads to SOTA performance. High AUC scores and low pose errors are preferred, thus, threshold = 0.15 is chosen, i.e., the patches with the descriptor similarities lower than 0.15 are recognized as neighbors.
@@ 5 Ablations
Dropout Layer. Table 7 shows the relative pose estimation performance on the top-$k$ images retrieved by the model with the searching radius and votings chosen from the best AUC, with or w/o global prefilter ([CLS] tokens), and dropout layer. Testing sets of MegaDepth are used. The CLS prefiltering improves the AUC scores. However, it increases the median pose errors marginally at the same time. As shown in the fifth row, data augmentation is essential in robustly learning the embeddings. Also, the last two rows in Table 7 show that the dropout layer improves the performance on MegaDepth and helps with generalization for pose estimation on other data and indoor localization.
@@
6 Conclusion
We introduce a novel Visual Place Recognition method, VOP, focusing on patch-level visual overlap prediction to address challenging viewpoint changes and occlusions. The proposed VOP predicts overlapping image patches by learning a patch-level descriptor and employing a robust voting mechanism and radius search in the descriptor space. Through extensive testing on the MegaDepth, PhotoTourism, ETH3D, and InLoc datasets, we demonstrate the effectiveness of the proposed VOP. On average, it achieves the highest accuracy, improving upon the state-of-the-art AnyLoc by two AUC points and reducing its median pose error by 25%. We believe this research marks a significant step forward in VPR, breaking the frame of image similarity prediction (through global or local features) to obtain a patch-level understanding of the query and database images.

### v3

@@ 4 Experimental Results
Our experiments include estimating two-view epipolar geometry and localizing query images from retrieved images. The VOP model is trained on 153 scenes from the MegaDepth dataset [33], which includes Internet-sourced images and depths reconstructed via Structure-from-Motion (SfM) and multi-view stereo techniques. Beyond two MegaDepth testing scenes, we evaluate the generalization ability of the proposed VOP on 11 scenes from PhotoTourism [53], 13 scenes from ETH3D [49], and the InLoc dataset [55]. We benchmark the proposed method against several competitors, including CNN-based retrieval methods like NetVLAD [3] and CosPlace [7], as well as the state-of-the-art (SOTA) retrieval method AnyLoc [30] and DINOv2-SALAD [29] utilizing the same DINOv2 backbone as ours. Besides those global-only methods, we compare VOP with the SOTA reranking-based methods: the ViT-based $R^{2}$ormer [68] and Patch-NetVLAD [23]. Also, we finetune CosPlace* on the same training data as ours to show the improved performance is not due to the different training data. In addition, we test the reranking methods [23] [68] on the shortlists created by DINOv2 [CLS] tokens, same as VOP, marked with $\dagger$ in the tables.
@@
Training Details. The models underwent training on the 153 MegaDepth scenes, employing contrastive loss to differentiate negative and positive patch pairs. Each scene contributes 150 positive and 150 negative image pairs to the training set, standardized to an image resolution of $224^{2}$ pixels. We use the pre-trained DINOv2 with a patch size of $14^{2}$ to extract $1024$-dimensional descriptors and reduce them to 256 dimensions by a fully-connected layer, i.e., the first layer of our training encoder head. Negative images, characterized by zero overlaps (e.g., sourced from distinct scenes), contrast with positive samples randomly selected from the pairs with 10% to 70% overlapping patches. Notably, the contrastive loss is predicated on patch-level negative/positive distinctions, i.e., cosine similarities, rather than on entire images. Following LightGlue [34], the patch-level supervision is generated from depth and visual features.The training was conducted in 30 epochs with a batch size of 64, employing the validation losses as the primary criterion for best checkpoint selection with a learning rate of $41e-$ 4.
@@
Evaluation Metrics. In the evaluation, we use the proposed weighted voting procedure with a radius search thresholded at the rounded median similarity over 100 random image samples. Note that the TF-IDF weights are used only when they are non-zero. Then, we assess the relative poses between the query and the top-$k$ retrieved images using robust estimation techniques. This involves deploying SuperPoint [14] and the state-of-the-art LightGlue [34, 48] for feature detection and matching, followed by OpenCV-RANSAC [20] running for 10K iterations to compute the relative poses. The performance is measured against ground truth (GT) poses by calculating the Area Under the Recall curve (AUC) thresholded at $10^{\circ}$ [65, 9], along with the median pose error. The RANSAC threshold is automatically selected from a set of pixel values $\{0.5,1,2,3\}$ following GlueFactory [34]. Also, we report the average number of inliers across the retrieved image pairs found by RANSAC. Note that the number of inliers is not an indicator of the accuracy of the estimated pose, as discussed and verified in MQNet [5] and FSNet [6]. In scenarios where $k>1$, for each query, we follow the practical approach and select only the image pair with the highest number of inliers. This comprehensive evaluation allows us to thoroughly assess the performance not just in retrieving relevant images but also in accurately estimating their relative poses.
@@
VOP 61.8 67.6 66.7 2.36 2.02 2.22 177.5 246.5 251.0
Table 1: Pose AUCs and median errors on MegaDepth [33] using top-$k$ images retrieved by different methods, with the best results in bold at each $k$ and the overall best underlined. results are tested by reranking images prefiltered by the [CLS] tokens from DINOv2. * shows results of fine-tuned CosPlace on Megadepth.
4.1 Pose Estimation on MegaDepth
Table 1 presents the results on the MegaDepth. The proposed method secures comparable scores for the AUC@10 criterion and exhibits the lowest median pose errors among the SOTA baselines. The underlined results mark the optimal configuration for relative pose accuracy, evidencing the superior performance of the proposed VOP approach when utilizing the top-5 images. This configuration outperforms the global-only retrieval baselines (the first seven rows), reinforcing the significance of selecting correct metrics and model configurations tailored to the specific demands of real-world downstream applications. Our method outperforms the recent reranking approaches $R^{2}$ormer [68] and Patch-NetVLAD [23] on MegaDepth when using their pretrained models. When using the same shortlist we use, their performance (with $\dagger$) is marginally better than ours. The fine-tuned CosPlace (CosPlace*) achieves better accuracy than the pretrained model, especially achieving good results on top-10 pairs. Next, we will show the generalization and robustness of the compared methods on unseen data.
AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
NetVLAD [3] 89.5 90.4 89.8 0.79 0.77 0.82 378.4 419.3 444.9
CosPlace [7] 88.3 89.8 89.8 0.82 0.80 0.80 352.3 421.6 434.9
CosPlace* [7] 43.6 79.7 85.9 38.3 0.79 0.76 162.6 279.0 360.5
DINOv2 [40] 87.2 89.9 89.6 0.82 0.83 0.86 365.2 415.0 436.7
AnyLoc [30] 88.9 90.2 90.2 0.77 0.83 0.80 373.2 425.1 438.0
SALAD [29] 89.7 90.2 90.0 0.68 0.76 0.78 385.9 411.2 442.0
P-NetVLAD [23] 86.7 88.4 90.2 0.81 0.83 0.81 333.3 412.4 466.8
P-NetVLAD [23] 80.5 88.1 89.4 0.84 0.77 0.74 291.7 341.5 372.2
$R^{2}$ormer[68] 50.2 72.0 84.4 34.0 1.58 0.84 137.0 318.3 401.5
$R^{2}$ormer[68] 51.1 83.0 87.4 13.4 0.79 0.86 114.7 272.5 314.7
VOP 89.7 90.5 90.4 0.69 0.70 0.70 341.1 399.5 430.7
Table 2: Relative pose AUC@10 and median errors on the unseen ETH3D dataset [49] on the top-$k$ retrieved images.
@@ 4.2 Generalization Experiments
VOP 62.5 66.8 65.2 2.63 2.12 2.44 224.5 250.5 283.2
Table 3: Relative pose AUC@10 and median errors on the unseen PhotoTourism dataset [53] on the top-$k$ retrieved images.
Method
@@
Avg. acc.(%) $\uparrow$ 70.1 71.6 56.4 70.6 73.2 72.3 66.3 65.6 74.8
Avg. err.($\downarrow$ 12.1 12.0 12.0 12.0 12.0 12.2 11.9 11.8 11.6
Table 5: Average accuracy (%) and median pose error () on the retrieved top-5 images on the Megadepth [33], ETH3D [49], PhotoTourism [53], and InLoc [55] datasets. InLoc is excluded from the median error as it is not available on the evaluation website.
Figure 5: The number of connected components (vertical axis, cc) is plotted against the index of the image pair (horizontal) on which RANSAC-based relative pose estimation runs. The left plot shows results for the top-1 database (DB) images paired with $0.4K$ query images, where the number of pairs equals the number of queries. The right plot shows results for the top-10 DB images with a termination criterion applied when all images are in a single component. Row “max $cc_{size}$” is the number of elements in the largest cc. Row “# cc” is the final number of connected components, while “$\text{idx}_{\text{last}}$” shows the index when the termination criterion was triggered in the right plot. Row “# skipped”, “# success”, and “# failure” show the numbers of skipped, succeeded, and failed RANSACs. All are in percentages.
InLoc. Additionally, we ran all methods on the InLoc dataset [55] (without fine-tuning) using hloc [48] for localization to show their generalization capabilities in the case of a large domain gap (in/outdoor). Following hloc [48], we retrieve top-40 database images using global retrieval methods and the reranking ones (including VOP) with a top-100 shortlist created by [CLS] tokens. Table 4 shows the proposed VOP achieves the second highest accuracy, showcasing its generalization abilities. Note that we only show results for P-NetVLAD and $R^{2}$ormer with DINOv2 [CLS] tokens as their original version fails on this dataset.
In summary, we show the average accuracy and median pose errors averaged over all tested datasets in Table 5. Note that InLoc is excluded from the median pose error, as the official website [57] does not report it. Overall, the proposed VOP leads to the highest accuracy and lowest pose errors. Fine-tuning CosPlace on Megadepth reduces its accuracy on average, demonstrating that our improved performance is not simply due to the different training data.
Figure 6: AUC@10 scores on the queries retrieving top-5 images from low-overlapping (horizontal axis) database.
@@ 4.3 Discussions
Pose Graph Construction. We run a pose graph construction experiment using the retrieved image pairs in the following way. All images are initialized to be connected components, each with a single element in it. We iterate through the queries and, for each, we add the query-to-database edge with the highest similarity/overlap to the reconstruction if and only if it connects two disconnected components by resulting in more than 10 inliers after running RANSAC. Thus, in each step, the number of connected components (cc) can only decrease by one or remain unchanged. Progressively, after all queries have been visited, we start again and process the edge with the second-highest similarity for each query. In the left (top-1) plot of Fig. 10, this is repeated until all edges have been visited; while termination criterion is used in the right (top-10), i.e., when all the images connected. Note that we repeat such a process for 1K times with shuffling the queries and plot their average.
The experiments were done on three test benchmarks: MegaDepth, PhotoTourism, and ETH3D. We show the averages by normalizing both horizontal and vertical axes. In the figures, the horizontal axis denotes the percentage of the currently processed image pairs (left - nothing processed so far, right - all pairs visited). The vertical axis is the percentage of connected components, i.e., the current number of cc divided by the initial. As shown in Fig. 10, our method results in the fewest connected components with the largest max cc size at top-1 case. In the right plot, $R^{2}$ormer and VOP are the only two methods that successfully connect all images in the pool, where VOP runs faster (more # skipped RANSACs). VOP results in the fewest RANSAC failures at top-1 case, and second-best at top-10 with 0.1% difference.
@@
See the ablations of data augmentation, dropout layer and radius threshold sensitivity in the Supp. Mat. It also includes qualitative results and more discussions about storage and speed, patching sizes, supervision options, etc.
5 Conclusion
We introduce a novel Visual Place Recognition method, VOP, focusing on patch-level visual overlap prediction to address challenging viewpoint changes and occlusion. The proposed VOP predicts overlapping image patches by learning a patch-level descriptor and employing a robust voting mechanism and radius search in the descriptor space. Through extensive testing on real-world indoor/outdoor datasets and low-overlap scenarios, we demonstrate the effectiveness of VOP. On average, it achieves the highest accuracy, improving upon the state-of-the-art AnyLoc by two AUC points and reducing its median pose error by 20%. Besides, VOP accelerates the pose graph construction process, tailored for geometric problems. We believe this research marks a significant step forward in VPR problems by patch-level understanding, and it tailors such methods for 3D reconstruction from unstructured image collections.
Acknowledgements. D. Barath was funded by the ETH Zurich Career Seed Award. J. Matas was supported by the Technology Agency of the Czech Republic, project No. SS05010008 and project No. SS73020004. T. Wei was funded by the Czech Science Foundation grant GA24-12697S and Research Center for Informatics (project CZ.02.1.01/0.0/0.0/16_019/0000765 funded by OP VVV).
Appendix
Appendix A Ablations
In this section, we will analyze certain components and parameters of the proposed method.
Radius Search Threshold. In the main paper, similar patches are selected by running radius search in the embedding space. We compute the median similarity as the radius search threshold over 100 random samples from the query and database images. Here, to understand how sensitive VOP is to the setting of this threshold, we show tuning results on the validation set of MegaDepth with different thresholds (horizontal axis) in Fig. 7. AUC scores and median errors are shown on the top-10 retrieved image pairs.
@@ Appendix A Ablations
Dropout Layer. Table 6 shows the relative pose estimation performance on the top-$k$ images retrieved by the model with or w/o global prefilter ([CLS] tokens), and dropout layer. Testing sets of MegaDepth are used, as in the main paper. The [CLS] prefiltering improves the AUC scores. However, it increases the median pose errors marginally at the same time. As shown in the fifth row, data augmentation is essential in robustly learning the embeddings. Also, the last two rows in Table 6 show that the dropout layer improves the performance on MegaDepth. As shown in the main paper, VOP generalizes well for pose estimation on other data and indoor localization.
@@
Figure 10: The number of connected components (vertical axis, cc) is plotted against the index of the image pair (horizontal) on which RANSAC-based relative pose estimation runs. The left plot shows results for the top-1 database (DB) images paired with $0.4K$ query images, where the number of pairs equals the number of queries. The right plot shows results for the top-10 DB images with a termination criterion applied when all images are in a single component. Row “max $cc_{size}$” is the number of elements in the largest cc. Row “# cc” is the final number of connected components, while “$\text{idx}_{\text{last}}$” shows the index when the termination criterion was triggered in the right plot. Row “# skipped”, “# success”, and “# failure” show the numbers of skipped, succeeded, and failed RANSACs. All are in percentages.
Appendix B Qualitative Results
Most VPR methods prioritize retrieving similar images, typically resulting in short baselines that are not suitable for reconstruction. These goals conflict: the most similar images often produce short baselines, making pose estimation unstable. We aim to move beyond traditional similarity metrics and design retrieval methods tailored for geometric challenges, such as selecting images suitable for pose estimation. We visualize three query examples in Fig. 8 with their top-1 retrieved images using different methods. VOP results in low pose errors as we find images with reasonable baselines for stable pose estimation.
Appendix C Discussions
Supervision. As discussed in the main paper, we build the supervision based on the ground-truth depth provided by Megadepth for training. However, some datasets do not provide depth information. Thus, we provide another option to build the supervision, i.e., patch-level positive and negative labels by matching the images using the SOTA dense feature matching method, RoMA [18]. The patch pairs containing more than $5$ correspondences are set as positive, and vice versa. As shown in Table 7, RoMA-based supervision works comparable to the model trained with depth supervision. Thus, we recommend using RoMA dense correspondences as a labeling option when fine-tuning on new data.
Supervision
depth-based RoMA-based
Avg. accuracy (%) $\uparrow$ 74.8 74.1
Avg. med. error ($\downarrow$ 11.6 11.7
Table 7: Average accuracy (%) and median pose error () on the retrieved top-5 images using different supervision on the test sets as the main paper. InLoc is excluded from the median error.
Method Backbone train data pre-filter
NetVLAD [3] ResNet-18 [25] Pitts30k [60] -
CosPlace [7] ResNet-101 [25] SF-XL [7] -
CosPlace*[7] ResNet-101 MegaDepth -
DINOv2 [40] ViT-G14 [17] LVD-142M [40] -
AnyLoc [30] DINOv2 - -
SALAD [29] DINOv2 GSV-Cities [1] -
P-NetVLAD [23] NetVLAD Pitts30k& MSLS [64] NetVLAD
[…]
$R^{2}$ormer [68] ViT-S [17] MSLS [64] $R^{2}$ormer
[…]
VOP DINOv2 MegaDepth DINOv2-CLS
Table 8: Backbones, training data, and the perfilter methods used for reranking shortlist generation are listed.
Fairness Comparison. As shown in Table 8, we show the backbones, training data and prefiltering methods (for reranking methods). We include fine-tuned results for CosPlace and results of the reranking methods tested with the same prefilter as ours.
Pose Graph construction. Similar to Fig. 5 shown in the main paper, here, we replace 1K random orders of queries by their predicted similarity/overlap scores. As shown in Fig. 10, VOP built a larger connected component on top-1 of the queries than the competitors, also with the most number of successful RANSACs. On the right plot (top-10), VOP iterates more queries than $R^{2}$ormer when ranked by the scores. However, the run-time is comparable as VOP skipped more times of RANSAC runs.
Image Patching. We investigated the number of patches to be used to split the images. The experiments were conducted on the testing scenes of MegaDepth, with all the images resized to $224^{2}$. From the trained VOP model with a patch size of $14^{2}$, we can extract 256 patch descriptors. Then, average pooling is applied to aggregate the patch descriptors to different patch sizes, such as $28^{2}$, $56^{2}$, $112^{2}$, and $224^{2}$. For example, patch size = $224^{2}$ will lead to a single patch of an image. The retrieval is done on the same prefiltered image list, similarly as in the main paper. As shown in Table 9, the aggregated patches e.g., patch size=$224^{2}$, perform worse than $14^{2}$ on MegaDepth pose estimation and could not generalize well on Inloc localization. This demonstrates that patch-level features can potentially improve estimated pose and other geometric problems.
MegaDepth DUC1 DUC2
$224^{2}$ 67.0 2.09 30.8 24.4
$112^{2}$ 67.0 2.09 38.4 38.9
$56^{2}$ 66.5 2.17 48.5 57.3
$28^{2}$ 65.9 2.29 59.1 72.5
$14^{2}$ 67.6 2.03 72.2 77.1
Table 9: Ablations on different patch sizes used in inference time. We show the AUC@10 scores and median pose errors of the top-5 retrieved images on MegaDepth [33], and the recall@, 0.5m on the top-40 of the localization data, Inloc [55] (DUC1, DUC2).
Storage & Query Speed. As we reduce the dimensionality of the DINOv2 features from 1024 to 256, the embeddings of all patches of each image need a total of 512 kB, while, for AnyLoc, the storage per image is 384 kB. While we require slightly more storage than AnyLoc [30], the difference is small. Compared to storing the local features in the reranking-based methods, VOP costs less. In addition, we compare the time of querying an image from the database of different sizes using AnyLoc or VOP in seconds. Prefiltering top 20 images by DINOv2 [CLS] token and running VOP for reranking to get top-1 out of 500 images cost 0.009 seconds, while 0.003 for AnyLoc. Querying top-1 from 5K images (prefiltered to 100), our method costs 0.03 seconds, while AnyLoc runs in 0.02s. Note that VOP needs much more time for radius search without prefiltering. We recommend using VOP as a reranking method in retrieval.